A person is jumping from a height h0 = 10 m with zero initial velocity and a cricket ball is thrown directly upward with an initial velocity v0 = 9.8 m/s at the same instant when time t = 0 s as shown in Fig. 1. Consider h = 11.5 m and i and j to be the unit vectors along the X and Y axis. Gravity is acting along the j direction and the magnitude of gravitational acceleration is given by g = 9.81 m/s2 . The red dot on the person represents his center of mass. (a) if the person and the cricket ball consist of a system, find the position and velocity of the center of mass of the system at time t = 0.5 s. (b) If it takes 0.01 seconds for the person to come to complete rest when he hits the ground, calculate the average force acting on the person. (c) Now consider when the person hits the ground, he bends his knees such that the center of mass of his body is lowered by 15 cm. If the maximum force that the bone of his legs can withstand is 50000 N, will he break his legs?


1
Expert's answer
2021-12-09T11:49:48-0500

(a) "x_c=\frac{m_1x_1+m_2x_2}{m_1+m_2}=\frac{60\cdot5+0.2\cdot20}{60+0.2}=5.05\ (m)"


"y_c=\frac{m_1\cdot(11.5-gt^2/2)+m_2\cdot(9.8t-gt^2/2)}{m_1+m_2}="


"=\frac{60\cdot(11.5-9.81\cdot0.5^2/2)+0.2\cdot(9.8\cdot0.5-9.81\cdot0.5^2/2)}{60+0.2}=10.25\ (m)"


"v_{cx}=\frac{m_1v_{1x}+m_2v_{2x}}{m_1+m_2}=\frac{m_1\cdot0+m_2\cdot0}{m_1+m_2}=0"


"v_{cy}=\frac{m_1\cdot(-gt)+m_2\cdot (9.8-gt)}{m_1+m_2}="


"=\frac{60\cdot(-9.81\cdot0.5)+0.2\cdot (9.8-9.81\cdot 0.5)}{60+0.2}=-4.87\ (m/s)"


"v_c=v_{cy}=-4.87\ (m/s)"



(b) "F=\frac{\Delta(mv)}{\Delta t}=m\cdot\frac{\Delta v}{\Delta t}=60\cdot\frac{\sqrt{2\cdot9.81\cdot10}-0}{0.01}=84\ 043\ (N)"


(c) using work-energy theorem, we get


"F_{floor}=mg(\frac{h}{\Delta h}+1)=60\cdot9.8\cdot(\frac{10}{0.15}+1)=39829\ (N)<50\ 000\ (N)"


The legs will be intact.